need help on this one! thanks
write a quadratic equation that has the solutions x=-5 and x=2

Answer by jai_kos(139)

Given the roots of the equation, we have x = -5 or x =2.
That is x =-5 or x =2
Get the numbers to the other side, we get
(x + 5) = 0 or (x -2) = 0
Or usually represent multiplication, we get
(x +5) * (x -2) = 0
Using the foil method, we multiply and get…
x^2 + 5x –x -10 = 0
x^2 + 4x -10 =0
The above represent the quadratic equation when x =-5 and x = 2